# Spatial-Network Treatment Effects: A Continuous Functional Approach

### v3

@@ Abstract
This paper develops a continuous functional framework for treatment effects that propagate through geographic space and economic networks. We derive a master equation governing propagation from three economic foundations—heterogeneous agent aggregation, market equilibrium, and cost minimization—establishing that the framework rests on fundamental principles rather than ad hoc specifications. A key result shows that the spatial-network interaction coefficient equals the mutual information between geographic and market coordinates. The Feynman-Kac representation decomposes effects into inherited and accumulated components along stochastic paths representing economic linkages. The framework nests the no-spillover case as a testable restriction. Monte Carlo simulations demonstrate that conventional estimators—two-way fixed effects, difference-in-differences, and generalized propensity score—exhibit 25–38% bias and severe undercoverage when spillovers exist, while our estimator maintains correct inference regardless of whether spillovers are present. Applying the framework to U.S. minimum wage policy, we reject the no-spillover null and find total effects at state borders four times larger than direct effects—conventional methods capture only one-quarter of policy impact. Structural estimates reveal spatial diffusion consistent with commuting-distance labor mobility, network diffusion consistent with quarterly supply chain adjustment, and significant spatial-network interaction reflecting geographic clustering of industries.
JEL Classification: C21, C23, J31, J38, R12
Keywords: treatment effects, spatial econometrics, network spillovers, continuous treatment
@@ 1.1 Contributions
The paper makes three main contributions spanning theory, methodology, and empirics.
Theoretical Contribution
The theoretical contribution derives a master equation governing treatment propagation from three independent economic foundations: heterogeneous agent aggregation following the tradition of Aiyagari (1994) and Huggett (1993), market equilibrium conditions connecting volatility to dynamic adjustment, and cost minimization principles underlying market equilibrium. The convergence of these derivations—each starting from different economic primitives but arriving at the same equation—establishes that the framework rests on fundamental economic principles rather than ad hoc specifications. The master equation
$\frac{\partial\tau}{\partial t}=\nu_{s}\nabla^{2}\tau+\nu_{n}\frac{\partial^{2}\tau}{\partial\alpha^{2}}-\kappa\tau+\lambda D_{s}D_{n}+S$ (1)
unifies spatial diffusion (coefficient $\nu_{s}$ reflecting labor mobility), network propagation ($\nu_{n}$ reflecting supply chain fluidity), market adjustment ($\kappa$ reflecting equilibration speed), spatial-network interaction ($\lambda$), and policy intervention ($S$) within a single analytical structure. Each parameter has a structural interpretation grounded in the underlying economic primitives.
A key theoretical result establishes that the spatial-network interaction coefficient $\lambda$ equals the mutual information between geographic and market coordinates. This information-theoretic characterization provides a parameter-free measure of channel complementarity grounded in the entropy concepts of Cover and Thomas (2006) rather than functional form assumptions. The Feynman-Kac representation connects the partial differential equation framework to economic intuition: treatment effects at any location reflect accumulated policy exposure along stochastic paths representing economic linkages—workers migrating, firms adjusting supply chains, prices equilibrating across connected markets.
Methodological Contribution
The methodological contribution demonstrates that the framework nests the no-spillover case as a testable restriction. When $\nu_{s}=\nu_{n}=0$, the master equation reduces to independent dynamics at each location, recovering standard continuous treatment estimation following Hirano and Imbens (2004) and Kennedy et al. (2017). This nested structure creates a one-sided risk profile: when spillovers are absent, the framework correctly detects this and produces estimates equivalent to conventional methods; when spillovers are present, the framework captures them while conventional methods fail.
We develop a comprehensive identification strategy combining three complementary approaches: spatial regression discontinuity exploiting sharp policy boundaries at state borders, network instrumental variables using predetermined supply chain connections, and entropy-based moment conditions leveraging theoretical predictions. Monte Carlo simulations with both binary treatment timing and continuous treatment intensity demonstrate that two-way fixed effects, heterogeneity-robust difference-in-differences, and generalized propensity score estimators exhibit 25–38% bias and 52–68% confidence interval coverage when spillovers of empirically relevant magnitude exist, while our estimator maintains correct inference across all configurations.
Empirical Contribution
The empirical contribution applies the framework to U.S. minimum wage policy, testing the no-spillover null and estimating structural propagation parameters. We reject the no-spillover null at $p<0.001$ and find that total effects at state borders are four times larger than direct effects alone—conventional methods capture only one-quarter of policy impact. Structural estimates reveal spatial diffusion consistent with commuting-distance labor mobility, network diffusion consistent with quarterly supply chain adjustment, and significant spatial-network interaction reflecting geographic clustering of industries.
@@ 1.2 Related Literature
This paper contributes to several literatures that have developed largely independently.
Spatial Econometrics and Spatial Inference
The spatial econometrics literature, surveyed by Anselin (2010) and formalized by LeSage and Pace (2009), models spillovers through spatial weight matrices capturing geographic proximity. This approach assumes spillovers decay with distance according to prespecified functional forms and cannot accommodate network connections that operate independent of geography. Conley (1999) develops spatial HAC inference but does not address network dependence.
@@
Recent work by Müller and Watson (2022) develops spatial correlation robust inference that remains valid under unknown forms of spatial dependence, providing confidence intervals with correct coverage regardless of the spatial correlation structure. Their approach addresses the critical problem that standard spatial HAC estimators require correct specification of spatial correlation decay. Müller and Watson (2024) further establishes conditions under which spatial data exhibit unit root behavior, leading to spurious regression phenomena analogous to time series contexts. These papers highlight fundamental inferential challenges in spatial settings that motivate careful attention to dependence structures. Our framework complements this literature by deriving the spatial dependence structure from economic primitives—the master equation implies specific correlation patterns determined by the diffusion and decay parameters—enabling counterfactual analysis that purely statistical approaches cannot provide.
Network Econometrics
The network econometrics literature, including Bramoullé et al. (2009) on peer effects identification and Jackson (2008) on network formation, addresses spillovers through social and economic connections. However, this literature typically treats network position as discrete (node membership) rather than continuous, ignoring geographic variation in network effects. Acemoglu et al. (2012) and Barrot and Sauvagnat (2016) document production network propagation but estimate reduced-form effects rather than structural propagation parameters. Our framework integrates network and spatial dimensions, models network position as continuous market coordinates, and recovers structural parameters governing propagation dynamics.
Treatment Effects Under Interference
The treatment effects literature has developed sophisticated methods for heterogeneity and selection, including generalized propensity scores for continuous treatments (Hirano and Imbens, 2004), doubly robust estimation (Kennedy et al., 2017), and heterogeneity-robust difference-in-differences (Callaway and Sant’Anna, 2021; Sun and Abraham, 2021; Goodman-Bacon, 2021). These advances address important problems but maintain SUTVA, ruling out spillovers by assumption.
A growing literature relaxes SUTVA to allow interference, including Hudgens and Halloran (2008) on partial interference and Athey et al. (2018) on experimental design under interference. However, this literature typically assumes known interference structure rather than estimating propagation dynamics from data. Our framework estimates the interference structure—the diffusion coefficients $(\nu_{s},\nu_{n})$ governing how treatment propagates—rather than assuming it known.
@@ 1.3 Outline
Section 2 develops the theoretical framework from microeconomic foundations, deriving the master equation from heterogeneous agent behavior, market equilibrium, and cost minimization, and establishing the Feynman-Kac representation and mutual information characterization. Section 3 establishes identification conditions, develops the no-spillover benchmark as a testable special case, presents comprehensive Monte Carlo evidence, and develops the GMM estimation framework with spatial-network HAC inference. Section 4 summarizes the empirical application to minimum wage spillovers. Section 5 concludes.
@@ 2.1 Coordinate System and Treatment Representation
Definition 1 (Spatial and Network Coordinates).
@@
Definition 2 (Source Term).
The source term $S(\mathbf{x},t,\alpha)$ represents the exogenous policy intervention entering the system. For minimum wage policy, $S$ equals the dollar amount by which the new minimum wage exceeds the market-clearing wage at each location-time-industry cell.
@@ 2.2 Derivation from Heterogeneous Agent Aggregation
Consider a continuum of heterogeneous agents indexed by type $\theta\in\Theta$ distributed over space and market positions. Each agent has idiosyncratic characteristics—productivity, preferences, constraints—captured by $\theta$. Agent $i$ of type $\theta$ at location $\mathbf{x}$ with market position $\alpha$ experiences wage:
$w_{i}(\mathbf{x},t,\alpha,\theta)=w_{0}(\mathbf{x},\alpha)+\tau(\mathbf{x},t,\alpha)+\varepsilon_{i}(\theta)$ (2)
where $w_{0}$ is the baseline wage, $\tau$ is the treatment effect to be determined, and $\varepsilon_{i}(\theta)$ captures idiosyncratic variation.
@@
where $(B_{t}^{s},B_{t}^{n})$ are independent Brownian motions representing location and occupational uncertainty. The drift terms $\mu_{s},\mu_{n}$ capture directed mobility—workers moving toward higher wages or better opportunities. The diffusion terms $\sigma_{s},\sigma_{n}$ capture randomness in mobility outcomes—search frictions, information imperfections, unexpected opportunities.
@@
Proposition 1 (Aggregation Result).
@@ 2.3 Derivation from Market Equilibrium
Connecting observed wage dynamics to treatment propagation: spatial wage volatility $\sigma_{s}^{2}$ implies spatial diffusion $\nu_{s}=\sigma_{s}^{2}/(2\kappa)$; network wage volatility $\sigma_{n}^{2}$ implies network diffusion $\nu_{n}=\sigma_{n}^{2}/(2\kappa)$. Markets with high wage volatility (active reallocation, frequent job changes) exhibit rapid spatial diffusion; markets with low volatility exhibit slow diffusion.
@@ 2.4 Derivation from Cost Minimization
Definition 3 (Adjustment Cost Functional).
@@
  • $\frac{1}{2}(\partial\tau/\partial t)^{2}$: Temporal adjustment costs. Rapidly changing wages are costly due to contracting frictions, menu costs, and coordination failures.
  • $\frac{\nu_{s}}{2}|\nabla\tau|^{2}$: Spatial gradient costs. Maintaining wage differentials across space is costly due to arbitrage pressure from mobile workers.
  • $\frac{\nu_{n}}{2}(\partial\tau/\partial\alpha)^{2}$: Network gradient costs. Maintaining wage differentials across market positions is costly due to substitution pressure in supply chains.
  • $\frac{\kappa}{2}\tau^{2}$: Level costs. Deviating from baseline wages is costly due to competitive pressure.
  • $S\tau$: Policy benefits. The intervention $S$ shifts the optimal wage structure.
Proposition 2 (Euler-Lagrange Equation).

### v4

# Treatment Effects with Correlated Spillovers: Bridging Discrete and Continuous Methods

@@ Abstract
This paper develops a continuous functional framework for treatment effects propagating through geographic space and economic networks. We derive a master equation from three independent economic foundations—heterogeneous agent aggregation, market equilibrium, and cost minimization—establishing that the framework rests on fundamental principles rather than ad hoc specifications. The framework nests conventional econometric models—autoregressive specifications, spatial autoregressive models, and network treatment effect models—as special cases, providing a bridge between discrete and continuous methods. A key theoretical result shows that the spatial-network interaction coefficient equals the mutual information between geographic and network coordinates, providing a parameter-free measure of channel complementarity. The Feynman-Kac representation characterizes treatment effects as accumulated policy exposure along stochastic paths representing economic linkages, connecting the continuous framework to event study methodology. The no-spillover case emerges as a testable restriction, creating a one-sided risk profile where correct inference is maintained regardless of whether spillovers exist. Monte Carlo simulations confirm that conventional estimators exhibit 25–38% bias when spillovers are present, while our estimator maintains correct inference across all configurations including the no-spillover case.
JEL Classification: C21, C23, C51
Keywords: treatment effects, spatial econometrics, network spillovers, continuous treatment, SUTVA
@@ 1.1 Contributions
The paper makes three main contributions. First, we derive a master equation governing treatment propagation from three independent economic foundations, establishing that the framework rests on fundamental principles rather than ad hoc specifications. A key theoretical result shows that the spatial-network interaction coefficient equals the mutual information between geographic and market coordinates, providing a parameter-free measure of channel complementarity. Second, we demonstrate that the framework nests the no-spillover case as a testable restriction, creating a one-sided risk profile where correct inference is maintained regardless of whether spillovers exist. Monte Carlo evidence confirms that conventional methods exhibit 25–38% bias when spillovers are present, while our framework maintains correct inference across all configurations. Third, applying the framework to U.S. minimum wage policy, we find total effects four times larger than direct effects alone, with structural parameters consistent with commuting-distance labor mobility and quarterly supply chain adjustment.
@@ 1.2 Related Literature
This paper contributes to several literatures that have developed largely independently, bridging spatial econometrics, network econometrics, and causal inference under interference.
1.2.0.1 Spatial Econometrics and Spatial Inference.
The spatial econometrics literature, surveyed by Anselin (2010) and formalized by LeSage and Pace (2009), models spillovers through spatial weight matrices capturing geographic proximity. This approach has proven successful for analyzing regional interdependencies and local economic shocks. However, it faces two fundamental limitations that motivate our approach. First, standard spatial models assume spillovers decay with distance according to prespecified functional forms (typically exponential or power-law kernels). While functional form misspecification can be tested, the structural content of these decay patterns remains opaque—estimates reveal correlations but not the underlying economic mechanisms. Second, spatial weight matrices capture only geographic proximity and cannot accommodate network connections that operate independent of location, such as supply chain relationships connecting distant firms or occupational networks spanning regions.
@@
Conley (1999) develops spatial HAC inference that accounts for unknown spatial correlation structures, providing a foundation for robust inference in spatial settings. Recent work by Müller and Watson (2022) extends this approach by developing spatial correlation robust inference that remains valid under unknown forms of spatial dependence, addressing the critical problem that standard spatial HAC estimators require correct specification of spatial correlation decay. Their confidence intervals maintain correct coverage regardless of the spatial correlation structure, providing stronger guarantees than conventional methods. Müller and Watson (2024) further establishes that spatial data can exhibit unit root behavior analogous to time series, leading to spurious regression phenomena even in cross-sectional spatial settings. These papers highlight fundamental inferential challenges that motivate our approach.
Our framework contributes to this literature by deriving the spatial dependence structure from economic primitives. The master equation implies specific correlation patterns determined by the diffusion coefficient $\nu_{s}$ and decay parameter $\kappa$, providing structural interpretations for the decay rate (labor mobility and commuting patterns) and adjustment speed (market equilibration). This structural foundation enables counterfactual analysis—predicting treatment effects under alternative policy designs or market structures—that purely statistical approaches to spatial correlation cannot provide. Moreover, by jointly modeling spatial and network channels, we address the limitation that geographic weight matrices miss economically important connections operating through supply chains, knowledge networks, or financial linkages.
1.2.0.2 Network Econometrics.
The network econometrics literature addresses spillovers through social and economic connections. Bramoullé et al. (2009) establish identification conditions for peer effects in social networks, showing that network structure provides variation separating direct effects from spillovers. Their reflection problem—distinguishing whether individual $i$’s outcome affects $j$ or vice versa—is a fundamental challenge when agents interact. Jackson (2008) provides a comprehensive treatment of network formation and the economic implications of network structure, establishing how network topology affects information diffusion, coordination, and strategic behavior.
Empirical applications have documented the quantitative importance of network spillovers across diverse settings. Acemoglu et al. (2012) show that production network structure substantially amplifies the impact of firm-level shocks on aggregate volatility, with the distribution of firm sizes and the topology of input-output linkages determining whether shocks dissipate or propagate system-wide. Barrot and Sauvagnat (2016) document that supply chain disruptions from natural disasters propagate through customer-supplier networks, affecting firms far removed geographically from the initial shock. These papers demonstrate that network position matters fundamentally for economic outcomes.
However, this literature typically treats network position as discrete—agents are nodes in a graph with discrete edges—rather than continuous, and ignores geographic variation in network effects. A California firm’s supply chain connections matter differently depending on whether suppliers are in neighboring Nevada or distant New York, yet discrete network models treat all edges symmetrically regardless of geographic proximity. Our framework addresses this by modeling market position as a continuous coordinate $\alpha$ capturing position in economic networks (production stage, skill level, risk profile) and jointly estimating how treatment effects propagate through both network and geographic dimensions. The interaction term $\lambda$ captures amplification when network connections concentrate geographically, a phenomenon discrete network models cannot accommodate.
Moreover, while existing network econometrics estimates reduced-form spillover effects (measuring correlations between connected agents), our framework recovers structural propagation parameters. The network diffusion coefficient $\nu_{n}$ measures how rapidly treatment effects transmit through supply chains or occupational networks, enabling counterfactual analysis of network restructuring or policy targeting based on network position. This structural approach complements the existing literature’s focus on identification and reduced-form measurement.
1.2.0.3 Treatment Effects Under Interference.
The treatment effects literature has developed sophisticated methods for heterogeneity and selection. Hirano and Imbens (2004) extend propensity score methods to continuous treatments, enabling estimation of dose-response functions without discretizing treatment intensity. Kennedy et al. (2017) develop doubly robust estimators that maintain consistency if either the outcome model or propensity score model is correctly specified, providing robustness to model misspecification. Recent advances in difference-in-differences address heterogeneous treatment effects with staggered adoption: Callaway and Sant’Anna (2021), Sun and Abraham (2021), and Goodman-Bacon (2021) develop estimators that remain unbiased when treatment effects vary across units and over time, resolving problems with conventional two-way fixed effects identified by Bertrand et al. (2004).
These methodological advances represent substantial progress in causal inference, but all maintain the stable unit treatment value assumption (SUTVA), which rules out spillovers by construction. Under SUTVA, unit $i$’s outcome depends only on $i$’s own treatment assignment, not on the treatment of other units. This assumption enables clean identification but fails in settings where economic connections generate spillovers—precisely the contexts where spatial and network effects matter most.
A growing literature relaxes SUTVA to allow interference. Hudgens and Halloran (2008) formalize partial interference, where units can be partitioned into clusters with interference within but not across clusters. Athey et al. (2018) develop experimental designs and exact tests for settings with network interference, providing finite-sample inference when randomization occurs on networks. Vázquez-Bare (2022) develop methods for spatial interference when treatment and control units are geographically close, showing how to construct valid estimators accounting for geographic spillovers.
However, this interference literature typically assumes the interference structure is known—which units affect which others and through what functional form—rather than estimating the propagation mechanism from data. For example, Hudgens and Halloran (2008) assume cluster assignments are known, and Athey et al. (2018) assume the network structure is observed and correctly specified. When the interference structure is misspecified, these methods can produce severely biased estimates.
Our contribution is to estimate the interference structure—the diffusion coefficients $(\nu_{s},\nu_{n})$ governing how treatment propagates through space and networks—rather than assuming it known. The master equation provides a structural model of propagation dynamics, with parameters interpretable in terms of labor mobility, supply chain adjustment, and market equilibration. The framework nests the no-spillover case as a testable restriction, enabling formal tests of whether interference exists before imposing interference assumptions. This approach complements the existing literature: when interference is known to exist (for example, in randomized saturation designs where researchers deliberately manipulate exposure intensity), our framework quantifies the propagation mechanism; when interference is uncertain, the framework tests whether it is empirically relevant.
1.2.0.4 Connections to Other Fields.
The continuous functional approach draws on mathematical techniques developed in physics and applied mathematics. The master equation parallels reaction-diffusion systems studied in chemical kinetics and biological pattern formation, where the Laplacian operator captures spatial spreading and reaction terms capture local dynamics. The Feynman-Kac representation originates in quantum mechanics and mathematical finance, providing a probabilistic interpretation of partial differential equation solutions through path integrals. These mathematical tools have proven powerful for analyzing complex systems in physics and biology; our contribution is adapting them to economic contexts with careful attention to economic foundations, identification challenges, and policy-relevant interpretations.
Anderson (1972) argues that complex systems exhibit emergent properties not reducible to individual components—aggregate behavior qualitatively different from micro-level mechanisms. The spatial-network interaction in our framework exemplifies this principle: when geographic and network coordinates are correlated (industries cluster geographically), the interaction term $\lambda$ captures amplification effects absent when channels operate independently. This emergence of interaction effects from the joint correlation structure mirrors Anderson’s insight about complex systems, while our information-theoretic characterization ($\lambda$ equals mutual information) provides a rigorous quantification.
@@ 1.3 Outline
Section 2 develops the theoretical framework from microeconomic foundations, deriving the master equation from heterogeneous agent behavior, market equilibrium, and cost minimization, and establishing the Feynman-Kac representation and mutual information characterization. Section 3 establishes identification conditions, develops the no-spillover benchmark as a testable special case, presents comprehensive Monte Carlo evidence demonstrating the one-sided risk profile property, and develops the GMM estimation framework with spatial-network HAC inference. Section 4 concludes.
@@ 2.1 Coordinate System and Treatment Representation
Definition 2.1 (Spatial and Network Coordinates).
@@
Definition 2.2 (Source Term).
The source term $S(\mathbf{x},t,\alpha)$ represents the exogenous policy intervention entering the system. For a generic policy, $S$ measures the intensity of direct policy exposure at each location-time-market cell.
@@ 2.2 Derivation from Heterogeneous Agent Aggregation
Consider a continuum of heterogeneous agents indexed by type $\theta\in\Theta$ distributed over space and market positions. Each agent has idiosyncratic characteristics—productivity, preferences, constraints—captured by $\theta$. Agent $i$ of type $\theta$ at location $\mathbf{x}$ with market position $\alpha$ experiences outcome:
$Y_{i}(\mathbf{x},t,\alpha,\theta)=Y_{0}(\mathbf{x},\alpha)+\tau(\mathbf{x},t,\alpha)+\varepsilon_{i}(\theta)$ (1)
where $Y_{0}$ is the baseline outcome, $\tau$ is the treatment effect to be determined, and $\varepsilon_{i}(\theta)$ captures idiosyncratic variation.
@@
where $(B_{t}^{s},B_{t}^{n})$ are independent Brownian motions representing location and network uncertainty. The drift terms $\mu_{s},\mu_{n}$ capture directed mobility—agents moving toward better opportunities. The diffusion terms $\sigma_{s},\sigma_{n}$ capture randomness in mobility outcomes—search frictions, information imperfections, unexpected opportunities.
@@
Proposition 2.1 (Aggregation Result).
@@ 2.3 Derivation from Market Equilibrium
Connecting observed outcome dynamics to treatment propagation: spatial outcome volatility $\sigma_{s}^{2}$ implies spatial diffusion $\nu_{s}=\sigma_{s}^{2}/(2\kappa)$; network outcome volatility $\sigma_{n}^{2}$ implies network diffusion $\nu_{n}=\sigma_{n}^{2}/(2\kappa)$. Markets with high outcome volatility (active reallocation, frequent adjustments) exhibit rapid spatial diffusion; markets with low volatility exhibit slow diffusion.
@@ 2.4 Derivation from Cost Minimization
Definition 2.3 (Adjustment Cost Functional).
@@
  • $\frac{1}{2}(\partial\tau/\partial t)^{2}$: Temporal adjustment costs. Rapidly changing outcomes are costly due to contracting frictions, menu costs, and coordination failures.
  • $\frac{\nu_{s}}{2}|\nabla\tau|^{2}$: Spatial gradient costs. Maintaining outcome differentials across space is costly due to arbitrage pressure from mobile agents.
  • $\frac{\nu_{n}}{2}(\partial\tau/\partial\alpha)^{2}$: Network gradient costs. Maintaining outcome differentials across market positions is costly due to substitution pressure in supply chains.
  • $\frac{\kappa}{2}\tau^{2}$: Level costs. Deviating from baseline outcomes is costly due to competitive pressure.
  • $S\tau$: Policy benefits. The intervention $S$ shifts the optimal outcome structure.
Proposition 2.2 (Euler-Lagrange Equation).